OpenStudy (anonymous):

Express the peice-wise function: \[ f(x) = \begin{cases} g(x)& x>a\\ h(x) &x

OpenStudy (anonymous):

The heaviside function is \[ H(t) = \begin{cases} 1&t>0\\ 0&t<0 \end{cases} \]

OpenStudy (anonymous):

If we put in \(t-a\) we get: \[ H(t-a) = \begin{cases} 1&t>a\\ 0&t<a \end{cases} \]

OpenStudy (anonymous):

Multiplying by \(g(t)\) we get: \[ g(t)H(t-a) = \begin{cases} g(t)&t>a\\ 0&t<a \end{cases} \]

OpenStudy (anonymous):

Another thing we could do is: \[ f(t)+[g(t)-f(t)]H(t-a) \]

OpenStudy (anonymous):

So one simple thing we could do is: \[ h(t)H(a-t) + g(t)H(t-a) \]

OpenStudy (anonymous):

Another thing we could do is: \[ f(t) =h(t)+[g(t)-h(t)]H(t-a) \]

OpenStudy (anonymous):

I'm producing a method of expressing peice-wise functions without using \(\large\{\) notation but instead by using \(H(t)\).

OpenStudy (anonymous):

Suppose we have \[ f(t) = \begin{cases} k(t)& t>a \\ m(t)& b<t<a\\ n(t)& t<b \end{cases} \]If we let \[ h(t) = \begin{cases} k(t)& t>a \\ m(t)& t<a \end{cases} \]Then we can say that: \[ f(t) = \begin{cases} h(t)& t>b \\ n(t)& t<b \end{cases} \]

OpenStudy (anonymous):

Thus, by induction, we could construct any piece-wise function, no matter how complicated, with Heaviside functions.